A builder is building a rectangular deck with dimensions of 16 feet by 30 feet. To ensure that the sides form 90 degree angles, what should each diagonal measure? (A). 30ft. (B). 34ft. (C). 46ft. (D). 16ft
Answer by nerdybill(7384)

apply Pythagorean theorem
Let d = diagonal
then
d^2 = 16^2 + 30^2
d^2 = 256 + 900
d^2 = 1156
d = 34 feet